Grade upto college level Electric Current

A particle is moving along X-axis whose acceleration is given by a = 3x-4 where x is the location of the particle.At t=0, the particle is at rest at x= 4/3 . The distance travelled by particle in 5 sec is

(a)zero (b)approx. 42m

(c)infinite (d)none of these

Ans ->(a)

Amit Saxena
12 Years agoGrade upto college level

1 Answer

mk chowdary
7 Years ago
At t=0, x=4/3 & initial velocity u=0
Acceleration a=3x-4
Therefore a=4(4/3)-4=0
As u=0 & a=0
I.e body remains at rest.
Hence distance travelled by the body in 5Sec is zero.